OpenStudy (anonymous):

Im not good at fractions. Help Please !? Add or subtract. 1/2 – 3/2

jigglypuff314 (jigglypuff314):

as long as the denominators are the same, you can add or subtract the numerators like you normally would so for example 1/3 - 3/3 = -2/3

OpenStudy (anonymous):

Ive came up with a few answers but they were wrong. Im confused. 2/2, 4/2 ….

jigglypuff314 (jigglypuff314):

well what is 1 - 3 = ?

OpenStudy (anonymous):

-2

jigglypuff314 (jigglypuff314):

yes, so you have -2/2 and remember that a number divided by itself = 1

OpenStudy (anonymous):

1/2-3/2=-2/2 ? thats what I came up with at first but, it was wrong. can a fraction be negative ?

jigglypuff314 (jigglypuff314):

that'd be my point from the second line :P a number divided by itself = 1 so for example 3/3 = 1 4/4 = 1 -4/4 = -1 etc

OpenStudy (jdoe0001):

\(\bf \cfrac{1}{2}-\cfrac{3}{2}=\cfrac{?}{{\color{brown}{ lcd?}}} \\ \quad \\ \to \cfrac{[({\color{brown}{ 2}}\div 2)\cdot 1]-[({\color{brown}{ 2}}\div 2)\cdot 3]}{{\color{brown}{ 2}}}\to \cfrac{1\cdot 1-1\cdot 3}{2}\to \cfrac{1-3}{2}\to \cfrac{-\cancel{ 2 }}{\cancel{ 2 }}\to -1\)

OpenStudy (anonymous):

THANK YOU SO MUCH !

OpenStudy (jdoe0001):

yw